# Enhanced Quantum Energy Teleportation using a 3-Qubit System

### v1

@@ II-B Deposit Energy (Alice)
Measurement of Alice’s $X$ operator is conducted, by which we can witness a state |+ket|$\rangle$| + ⟩ or |ket|$\rangle$| - ⟩. Bob’s energy is certainly not affected by this operation since $[X_{0},V_{0,1}]=[X_{0},H_{1}]=0$. Using $[P_{0}(\mu_{0}),V_{0,1}]=0$ and $\langle+|Z|+\rangle=\langle-|Z|-\rangle=0$, Alice’s mean energy to deposit can be calculated as,
@@ II-C Deposit Energy (Charlie)
Charlie’s $X$ operator is measured, by which he obtains a state |+ket|$\rangle$| + ⟩ or |ket|$\rangle$| - ⟩. His operation also does not affect Bob’s energy since $[X_{1},V_{0,2}]=[X_{0},H_{2}]=0$. Using $[P_{0}(\mu),V_{0,2}]=0$ and $\langle+|Z|+\rangle=\langle-|Z|-\rangle=0$, we find that Charlie’s mean energy to deposit is
@@
Step 2: Receive Energy
As soon as Alice observes $\mu_{0}\in\{-1,1\}$ and Chalie observes $\mu_{1}\in\{-1,1\}$ , They communicates their result to Bob who operates $U_{B}(\mu_{0,1})$ to his qubit and measures his energy. Here $U_{B}(\mu_{0,1})$ is obtained as
$U_{1}(\mu_{0,1})=\cos\phi I-i\mu\sin\phi Y_{1}=R_{Y}(2\phi),$ (30)
@@
$\rho_{\text{QET}}=\sum_{\mu\in\{-1,1\}}U_{1}(\mu_{0,1})P_{1}(\mu_{1})P_{0}(\mu_{0})|g\rangle\langle g|P_{0}(\mu_{0})P_{1}(\mu_{1})U_{1}^{\dagger}(\mu_{0,1})$ (31)
@@
$\langle E_{2}\rangle=\text{Tr}[\rho_{\text{QET}}(H_{2}+V_{0,1}+V_{0,2})]=\text{Tr}[\rho_{\text{QET}}H_{\text{tot}}]-(\langle E_{0}\rangle+\langle E_{1}\rangle),$ (32)
@@
Here is a little bit explanation of quantum circuits for the extended QET protocol. Similar to minimal QET model $V$ and $H_{2}$ do not commute, measurement of those has to be done separately. That means, Bob measures $V_{0,2}$,$V_{0,1}$ and $H_{2}$ independently and obtains evaluated $V_{0,2}$,$V_{0,1}$ and $\langle H_{2}\rangle$ statistically. We witness, $V_{0,2}$,$V_{1,2}$ to be always negative and $\langle H_{2}\rangle$ is always positive. Therefore it is sufficient for Bob to measure only $V_{0,2}$,$V_{1,2}$ to receive energy with our extended QET protocol.
@@
$\langle E_{2}\rangle=\sum_{\mu\in\{-1,1\}}\langle g|P_{0}(\mu_{0})P_{1}(\mu_{1})U_{1}^{\dagger}(\mu)(H_{2}+V_{0,1}+V_{0,2})U_{1}(\mu)P_{1}(\mu_{1})P_{0}(\mu_{0})|g\rangle.$ (33)
@@ IV Measurements and Quantum Gates
$\Lambda(U)=|0\rangle|0\rangle\langle 0|\langle 0|\otimes I+|1\rangle|1\rangle\langle 1|\langle 1|\otimes U$ (36)
@@
$CCNOT(R_{Y}(\theta_{1}\otimes I\otimes I)|000\rangle=0|000\rangle+a|001\rangle+b|010\rangle+0|011\rangle+c|100\rangle+0|101\rangle+0|110\rangle+d|111\rangle.$ (37)
$CNOT(R_{Y}(I\otimes\theta_{2}\otimes I)|000\rangle=0|000\rangle+a|001\rangle+b|010\rangle+0|011\rangle+c|100\rangle+0|101\rangle+0|110\rangle+d|111\rangle.$ (38)
@@ V Some details of the model
In order to comprehend the non-triviality of the QET protocol, it is crucial to acknowledge that regardless of the specific unitary operation W1 applied to Bob’s qubit following Alice’s measurement, it is impossible to extract any energy. This can be verified by the equation
$\text{Tr}[\rho_{W}H_{tot}]-\langle E_{0}\rangle=\langle g|W_{1}^{\dagger}H_{tot}W_{1}|g\rangle\geq 0,$ (41)
@@
$\rho_{W}=W_{1}^{\dagger}\left(\sum_{\lambda=\pm 1}P_{1}(\mu_{1})P_{0}(\mu_{0})|g\rangle\langle g|P_{0}(\mu_{0})P_{1}(\mu_{1})\right)W_{1}.$ (22)

### v2

@@ II-B Deposit Energy (Alice)
Measurement of Alice’s $X$ operator is conducted, by which we can witness a state |+ket|$\rangle$| + ⟩ or |ket|$\rangle$| - ⟩. Bob’s energy is certainly not affected by this operation since $[X_{0},V_{0,2}]=[X_{0},H_{0}]=0$. Using $[P_{0}(\mu_{0}),V_{0,1}]=0$ and $\langle+|Z|+\rangle=\langle-|Z|-\rangle=0$, Alice’s mean energy to deposit can be calculated as,
@@ II-C Deposit Energy (Charlie)
Charlie’s $X$ operator is measured, by which he obtains a state |+ket|$\rangle$| + ⟩ or |ket|$\rangle$| - ⟩. His operation also does not affect Bob’s energy since $[X_{1},V_{0,2}]=[X_{1},H_{2}]=0$. Using $[P_{0}(\mu),V_{0,2}]=0$ and $\langle+|Z|+\rangle=\langle-|Z|-\rangle=0$, we find that Charlie’s mean energy to deposit is
@@
Step 2: Receive Energy (Bob)
As soon as Alice observes $\mu_{0}\in\{-1,1\}$ and Chalie observes $\mu_{1}\in\{-1,1\}$ , They communicates their result to Bob who operates $U_{B}(\mu_{0},\mu_{1})$ to his qubit and measures his energy. Here $U_{B}(\mu_{0},\mu_{1})$ is obtained as
$U_{1}(\mu_{0},\mu_{1})=\cos\phi I-i\mu\sin\phi Y_{1}=R_{Y}(2\phi),$ (30)
@@
$\rho_{\text{QET}}=\sum_{\mu\in\{-1,1\}}U_{1}(\mu_{0},\mu_{1})P_{1}(\mu_{1})P_{0}(\mu_{0})|g\rangle\langle g|P_{0}(\mu_{0})P_{1}(\mu_{1})U_{1}^{\dagger}(\mu_{0},\mu_{1})$ (31)
@@
$\langle E_{2}\rangle=\text{Tr}[\rho_{\text{QET}}(H_{2}+V_{0,2}+V_{1,2})]=\text{Tr}[\rho_{\text{QET}}H_{\text{tot}}]-(\langle E_{0}\rangle+\langle E_{1}\rangle),$ (32)
@@
Here is a little bit explanation of quantum circuits for the extended QET protocol. Similar to minimal QET model $V$ and $H_{2}$ do not commute, measurement of those has to be done separately. That means, Bob measures $V_{0,2}$,$V_{1,2}$ and $H_{2}$ independently and obtains evaluated $V_{0,2}$,$V_{1,2}$ and $\langle H_{2}\rangle$ statistically. We witness, $V_{0,2}$,$V_{1,2}$ to be always negative and $\langle H_{2}\rangle$ is always positive. Therefore it is sufficient for Bob to measure only $V_{0,2}$,$V_{1,2}$ to receive energy with our extended QET protocol.
@@
$\langle E_{2}\rangle=\sum_{\mu\in\{-1,1\}}\langle g|P_{0}(\mu_{0})P_{1}(\mu_{1})U_{1}^{\dagger}(\mu)(H_{2}+V_{0,2}+V_{1,2})U_{1}(\mu)P_{1}(\mu_{1})P_{0}(\mu_{0})|g\rangle.$ (33)
@@ IV Measurements and Quantum Gates
$\Lambda(U)=|00\rangle\langle 00|\otimes I+|11\rangle\langle 11|\otimes U$ (36)
@@
$\begin{array}[]{l}\left[\left(I\otimes\left(|0\rangle\langle 0|\otimes I+|1\rangle\langle 1|\otimes X\right)\right)\cdot\left(I\otimes R_{y}(\theta)\otimes I\right)\cdot\left(X\otimes I\otimes I\right)\cdot\left(|0\rangle\langle 0|\otimes I\otimes I+|1\rangle\langle 1|\otimes I\otimes X\right)\cdot\right.\\ \left.\left(|0\rangle\langle 0|\otimes I\otimes I+|1\rangle\langle 1|\otimes X\otimes I\right)\cdot\left(R_{y}(\theta)\otimes I\otimes I\right)\cdot\left(X\otimes I\otimes I\right)\right]\cdot|000\rangle\\ =0|000\rangle+a|001\rangle+b|010\rangle+0|011\rangle+c|100\rangle+0|101\rangle+0|110\rangle+d|111\rangle.\end{array}$ (37)
@@ V Some details of the model
In order to comprehend the non-triviality of the QET protocol, it is crucial to acknowledge that regardless of the specific unitary operation W1 & W2 applied to Bob’s qubit following Alice’s and Charlie’s measurement, it is impossible to extract any energy. This can be verified by the equation
$\text{Tr}[\rho_{W}H_{tot}]-(\langle E_{0}\rangle+\langle E_{1}\rangle)=\langle g|W_{2}^{\dagger}W_{1}^{\dagger}H_{tot}W_{1}W_{2}|g\rangle\geq 0,$ (40)
@@
$\rho_{W}=W_{2}^{\dagger}W_{1}^{\dagger}\left(\sum_{\lambda=\pm 1}P_{1}(\mu_{1})P_{0}(\mu_{0})|g\rangle\langle g|P_{0}(\mu_{0})P_{1}(\mu_{1})\right)W_{1}W2.$ (22)